OpenStudy (anonymous):

how would you solve sin2x=0.5?

OpenStudy (anonymous):

sin(2x)=.5 2x=arcsin(.5) x = arcsin(.5)/2

OpenStudy (anonymous):

No but without a calculator?

OpenStudy (anonymous):

That is without a calculator.

OpenStudy (anonymous):

but how would you find the arcsin of .5/2?

OpenStudy (anonymous):

the actual number... i would use newton approximation

OpenStudy (anonymous):

lol. im only in precal -_- sorry i dont know about that

OpenStudy (amistre64):

sin(30) = 1/2 30 = 2x x = 60

OpenStudy (amistre64):

x also = 120

OpenStudy (amistre64):

....make that 15 lol

OpenStudy (amistre64):

and 180-15

OpenStudy (anonymous):

sin(x)=1/2 x = pi/6 sin(2x)=.5 x = pi/3

OpenStudy (anonymous):

pi/12, sorry

OpenStudy (anonymous):

wait wait wait hold on. Wouldnt x=15. Since sin of 30=.5? And sin of 2(15)=.5?

OpenStudy (anonymous):

In degrees it would be 15, in radians, pi/12

OpenStudy (anonymous):

ahhhhhh. oh ok. See now i know what im doing. bcuz at first i didnt even understand what the problem was asking me to find out.

OpenStudy (anonymous):

Thank you thank you. :)